# Enhancement of superconductivity outside Abrikosov vortex core in a tightly bound Cooper pair superconductor

### v3

@@ II Feynman’s phenomenological theory
where $\mathcal{L}$ is the Lagrangian density (hereafter called the Lagrangian) and $q$ is a shorthand for a full set of dynamical field quantities describing the state of the system. They include the real and imaginary parts of $\Psi$ (or equivalently $\Psi$ and $\Psi^{*}$), the scalar potential and components of the vector potential.
@@
It is instructive to start with the case of a superfluid of neutral particles of mass $m$. The particles may be ”elementary” bosons (such as $^{4}He$ atoms) or ”composite” bosons made of pairs of bound neutral spin-1/2 fermions. This system is described by the Lagrangian
@@
$V(n)=\frac{1}{2}\frac{ms^{2}}{n_{0}}(n-n_{0})^{2},~~~s^{2}=\frac{n_{0}}{m}V^{\prime\prime}(n_{0})>0$ (6)
where the parameter $s(n_{0})$ is the velocity of sound. We can now introduce the coherence length
$\xi=\frac{\hbar}{\sqrt{2}ms}$ (7)
@@ III Stationary field equations
$\frac{m\mathbf{v}^{2}}{2}+e\phi+\frac{ms^{2}}{n_{0}}(n-n_{0})-\frac{\hbar^{2}}{2m}\frac{\nabla^{2}\sqrt{n}}{\sqrt{n}}=0.$ (20)
@@ IV Abrikosov flux line
$\varepsilon=\frac{\mathcal{E}}{L_{z}}$ $=$ $\int\Big{\{}\frac{\hbar^{2}}{2m}(\nabla\sqrt{n})^{2}+\frac{ms^{2}}{2n_{0}}(n-n_{0})^{2}$ (23)
@@
$\mathbf{A}$ $=$ $\frac{\hbar cl}{e}\frac{1-F(\rho)}{\rho}\mathbf{e}_{\varphi},\phi=\frac{ms^{2}}{e}w(\rho).$ (27)
@@
$\lambda^{2}=\frac{mc^{2}}{4\pi n_{0}e^{2}},~~~\delta^{2}=\frac{ms^{2}}{4\pi n_{0}e^{2}},$ (30)
@@ IV.1 Field configuration
$\kappa\simeq\frac{s}{\alpha^{2}c},~~~\gamma\simeq\frac{s^{2}}{\alpha^{2}c^{2}}$ (34)
where $\alpha=e^{2}/\hbar c\approx 1/137$ is the fine structure constant. Assuming $c/s\approx 10^{2}$ (i.e. the speed of sound $s$ has the order of magnitude of the Fermi velocity in a normal metal), we discover that $\kappa\simeq 10^{2}$ while $\gamma\simeq 1$. Restated in terms of the original length scales of the problem, this implies that the London penetration depth $\lambda$ is significantly larger than the coherence length, $\lambda\gg\xi$. In the classification of Ginzburg and Landau and Abrikosov LL9 ; Abrikosov this is a strongly type-II superconductor.
@@ V Conclusions
To summarize, we demonstrated that at zero temperature electrostatic effects in tightly bound Copper pair superconductors qualitatively modify the structure of the Abrikosov flux line. The latter is now characterized by profiles of the magnetic and electric fields and the density of superconductive electrons. The latter exhibits a maximum that exceeds the bulk value of the superconductive density. While superconductivity is still degraded within the (smaller) vortex core, it is enhanced outside of it. The underlying physical mechanism behind this behavior is electrostatic screening that enforces global neutrality. We also demonstrated that in the limit of strong electrostatic screening profiles of the magnetic and electric fields attain a universal form that has its origin in quantization of the magnetic flux and existence of equilibrium bulk density of superconductive electrons.
@@
We hope experimental efforts will be undertaken to determine the detailed structure of individual Abrikosov flux lines and compare with our predictions. The most promising experimental technique appears to be the scanning tunneling microscopy (STM) STM1 ; STM2 . What is measured is the local density of states (LDOS) at different energies, which reflects the suppression of the superconducting gap (i.e. the density of superconductive electrons) as the vortex axis is approached. STM’s atomic scale spatial resolution makes is optimally suited to carry out detailed measurements of the profile of the density of superconductive electrons. Having determined the superconductive density, the profiles of the electric and magnetic fields can be inferred from the Poisson equation (32) and Ampère law (33), respectively. Suitable materials include both conventional superconductors such as $NbSe_{2}$, $Pb$ and copper-based superconductors of the Bismuth Strontium Calcium Copper oxide family.

### v4

# Enhancement of superconductivity outside an Abrikosov vortex core in a tightly bound Cooper pair superconductor

@@ II Feynman’s phenomenological theory
where $\mathcal{L}$ is the Lagrangian density (hereafter called the Lagrangian) and $q$ is a shorthand for a full set of dynamical field quantities describing the state of the system. They include the real and imaginary parts of the wave function $\Psi$ (or equivalently $\Psi$ and $\Psi^{*}$), the scalar potential and components of the vector potential.
@@
It is instructive to start with the case of a superfluid of neutral particles of mass $m$. The particles may be ”elementary” bosons (such as 4He atoms) or ”composite” bosons made of pairs of bound neutral spin-1/2 fermions. This system is described by the Lagrangian
@@
$V(n)=\frac{1}{2}\frac{mu^{2}}{n_{0}}(n-n_{0})^{2},~~~u^{2}=\frac{n_{0}}{m}V^{\prime\prime}(n_{0})>0$ (6)
where the parameter $u(n_{0})$ is the velocity of sound. We can now introduce the coherence length
$\xi=\frac{\hbar}{\sqrt{2}mu}$ (7)
@@ III Stationary field equations
$\frac{m\mathbf{v}^{2}}{2}+e\phi+\frac{mu^{2}}{n_{0}}(n-n_{0})-\frac{\hbar^{2}}{2m}\frac{\nabla^{2}\sqrt{n}}{\sqrt{n}}=0.$ (20)
@@ IV Abrikosov flux line
$\varepsilon=\frac{\mathcal{E}}{L_{z}}$ $=$ $\int\Big\{\frac{\hbar^{2}}{2m}(\nabla\sqrt{n})^{2}+\frac{mu^{2}}{2n_{0}}(n-n_{0})^{2}$ (23)
@@
$\mathbf{A}$ $=$ $\frac{\hbar cl}{e}\frac{1-F(\rho)}{\rho}\mathbf{e}_{\varphi},\phi=\frac{mu^{2}}{e}w(\rho).$ (27)
@@
$\lambda^{2}=\frac{mc^{2}}{4\pi n_{0}e^{2}},~~~\delta^{2}=\frac{mu^{2}}{4\pi n_{0}e^{2}},$ (30)
@@ IV.1 Field configuration
$\kappa\simeq\frac{u}{\alpha^{2}c},~~~\gamma\simeq\frac{u^{2}}{\alpha^{2}c^{2}}$ (34)
where $\alpha=(e/2)^{2}/\hbar c\approx 1/137$ is the fine structure constant. Assuming $c/u\approx 10^{2}$ (i.e. the speed of sound $u$ has the order of magnitude of the Fermi velocity in a normal metal), we discover that $\kappa\simeq 10^{2}$ while $\gamma\simeq 1$. Restated in terms of the original length scales of the problem, this implies that the London penetration depth $\lambda$ is significantly larger than the coherence length, $\lambda\gg\xi$. In the classification of Ginzburg and Landau and Abrikosov LL9 ; Abrikosov this is a strongly type-II superconductor.
@@ V Conclusions
To summarize, we demonstrated that at zero temperature electrostatic effects in tightly bound Cooper pair superconductors qualitatively modify the structure of the Abrikosov flux line. The latter is now characterized by profiles of the magnetic and electric fields and the density of superconductive electrons. The latter exhibits a maximum that exceeds the bulk value of the superconductive density. While superconductivity is still degraded within the (smaller) vortex core, it is enhanced outside of it. The underlying physical mechanism behind this behavior is electrostatic screening that enforces global neutrality. We also demonstrated that in the limit of strong electrostatic screening profiles of the magnetic and electric fields attain a universal form that has its origin in quantization of the magnetic flux and existence of equilibrium bulk density of superconductive electrons.
@@
We hope experimental efforts will be undertaken to determine the detailed structure of individual Abrikosov flux lines and compare with our predictions. The most promising experimental technique appears to be the scanning tunneling microscopy (STM) STM1 ; STM2 . What is measured is the local density of states (LDOS) at different energies, which reflects the suppression of the superconducting gap (i.e. the density of superconductive electrons) as the vortex axis is approached. STM’s atomic scale spatial resolution makes is optimally suited to carry out detailed measurements of the profile of the density of superconductive electrons. Having determined the superconductive density, the profiles of the electric and magnetic fields can be inferred from the Poisson equation (32) and Ampère law (33), respectively.
Suitable materials include both conventional superconductors like NbSe2, Pb, and copper-based superconductors belonging to the Bismuth Strontium Calcium Copper oxide family. For instance, since NbSe2 is a good conductor, electrostatic screening is exceptionally strong and occurs over a very short distance, approximately $\delta\simeq 0.1~nm$. Analysis and images of the original observations STM1 are consistent with a coherence length $\xi\simeq 10~nm$. Consequently, the dimensionless Debye screening length $\gamma=\delta/\xi$ (32) is of the order $0.01$. Our theory then predicts that the density of superconductive electrons outside the vortex core is about $10\%$ higher than its bulk counterpart.